# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ V-B Knowledge Augmentation
Before providing specific diagnoses and prescriptions, physicians typically inquire about patients’ medical and allergy histories. However, the capabilities of medical LLMs rely mainly on training data that do not contain information about the currently attending patients. Consequently, proposed diagnoses and prescriptions may not be applicable to individual patients. Therefore, a significant challenge is enabling medical LLMs to provide personalized services to patients or doctors using real-time information. Retrieval-Augmented Generation (RAG) [142] presents a feasible solution. This method first retrieves content relevant to the user query from external real-time sources (e.g., medical news, personal electronic medical records). It then integrates the user query with the retrieved content into a coherent prompt, and the model subsequently responds based on its own knowledge and the prompt. Several medical LLMs work have been developed to enhance responses using external knowledge, thereby improving performance and mitigating hallucination problems. For instance, ChatDoctor [25] employs online resources and custom offline medical databases as an external knowledge base to enhance accuracy in medical tasks through RAG. And JMLR [51] trains medical LLMs together with RAG to improve the model’s ability to process medical knowledge while reducing computational resources.
@@
However, existing medical LLMs do not fully leverage the capabilities of RAG. The description or medical terminology provided by the user’s limited level of medical knowledge may not accurately recall useful content. Therefore, research on keyword analysis and rewriting of user queries is necessary to improve knowledge retrieval accuracy and better utilize the model’s medical knowledge. Additionally, healthcare organizations typically maintain electronic medical records for their patients. Investigating how to use this information to assist medical LLMs in providing more personalized services to patients is a valuable research direction.
V-C Evaluation Platform for Medical Large Language Models
@@
Currently, several benchmarks or platforms exist for evaluating general LLMs, such as GLUE [143], SuperGLUE [144], and CLUE [145]. These platforms provide consistent evaluation criteria and standardized training and testing sets, facilitating fair comparisons and advancements among LLMs. Existing medical LLMs are evaluated using inconsistent datasets and ambiguous evaluation tasks, resulting in comparisons and improvements between models becoming difficult. Therefore, it is particularly necessary to establish a unified evaluation platform focused on medical LLMs to foster the advancement of this field.
@@
Establishing a unified evaluation platform for medical LLMs necessitates attention to several critical considerations. Firstly, the platform should incorporate benchmarks across various medical subdomains to ensure comprehensive evaluation of model performance across diverse areas. Secondly, the evaluation tasks should be clearly defined, encompassing medical QA, case diagnosis, medical reasoning, treatment recommendation, and other pertinent aspects to accurately reflect the model’s practical applicability. Lastly, the evaluation platform should encompass a diverse range of metrics and dimensions, including traditional measures such as accuracy and BLEU, alongside considerations of dimensions such as proactive, helpful, and safe.
VI Conclusion
@@
This survey provides a guide on training customized medical LLMs based on general LLMs. Customized medical LLMs can meet the personalized demands of individuals and healthcare organizations. We detail the training process at three levels: data, methodology, and evaluation. We analyze the data level in terms of corpus sources and processing methods, the methodology level in terms of training paradigms, and the evaluation level from both machine and human perspectives. For each level, we offer recommendations and guidance. Finally, we present the challenges and future research directions for each of these three levels.
Acknowledgments

### v2

@@ 5.3 Personalized Diagnosis
Before providing specific diagnoses and prescriptions, physicians typically inquire about patients’ medical and allergy histories. However, the capabilities of medical LLMs rely mainly on training data that do not contain information about the currently attending patients. Consequently, proposed diagnoses and prescriptions may not be applicable to individual patients. Therefore, a significant challenge is enabling medical LLMs to provide personalized services to patients or doctors using real-time information. Retrieval-Augmented Generation (RAG) [85] presents a feasible solution. This method first retrieves content relevant to the user query from external real-time sources (e.g., medical news, personal electronic medical records). It then integrates the user query with the retrieved content into a coherent prompt, and the model subsequently responds based on its own knowledge and the prompt [46]. Several medical LLM works have been developed to enhance responses using external knowledge, thereby improving performance and mitigating hallucination problems. For instance, ChatDoctor [91] employs online resources and custom offline medical databases as an external knowledge base to enhance accuracy in medical tasks through RAG. Moreover, JMLR [153] trains medical LLMs together with RAG to improve the model’s ability to process medical knowledge while reducing computational resources.
@@
However, existing medical LLMs do not fully leverage the capabilities of RAG. The description or medical terminology provided by the user’s limited medical knowledge may not accurately recall helpful content. Therefore, researchers may consider how to enrich or rewrite patient questions based on individual cases to improve information retrieval accuracy. Additionally, healthcare organizations typically maintain electronic medical records for their patients. Researchers can consider optimize the information retrieval embedding model for electronic health records to improve the accuracy and reduce the noise of retrieved information, thereby empowering medical large language models to enhance personalized diagnosis.
5.4 Unified Evaluation Platform
@@
Currently, several benchmarks or platforms exist for evaluating general LLMs, such as GLUE [149], SuperGLUE [148], CLUE [173], and PromptCBLUE [201]. These platforms provide consistent evaluation criteria and standardized training and testing sets, facilitating fair comparisons and advancements among LLMs. However, existing medical LLMs are evaluated using inconsistent datasets and ambiguous evaluation tasks, resulting in comparisons and improvements between models becoming difficult. Therefore, it is particularly necessary to establish a unified evaluation platform focused on medical LLMs to foster the advancement of this field.
@@
Establishing a unified evaluation platform for medical LLMs necessitates attention to several critical considerations. Firstly, the platform should incorporate benchmarks across various medical subdomains to ensure comprehensive evaluations of model performance across diverse areas. Secondly, the evaluation tasks should be clearly defined, encompassing medical QA, case diagnosis, medical reasoning, treatment recommendation, and other pertinent aspects to accurately reflect the model’s practical applicability. Lastly, the evaluation platform should encompass various metrics and dimensions, including traditional measures such as accuracy and BLEU, alongside considerations of dimensions such as proactive, helpful, and safe.
6 Conclusion
@@
This survey provides a guide on training customized medical LLMs based on general LLMs. We detail the training process at three levels: data, methodology, and evaluation. Specifically, we analyze the data level regarding corpus sources and processing methods, the methodology level in terms of training paradigms, and the evaluation level from both machine and human perspectives. For each level, we provide corresponding recommendations for organizations that desire to develop a customized medical LLM. Finally, we present the challenges and future research directions for these three levels.
[…]